# A Nonabelian Hodge Correspondence for Principal Bundles in Positive Characteristic

### v1

@@ 4.4. Stability Condition
Given $(\mathscr{E},\theta)\in{\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$, Theorem 4.12 shows that there is a logarithmic flat $G$-bundle $(\mathscr{V},\nabla)$ in ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$. Then, the proof is similar to that of Theorem 3.6 and we only have to check that there is a one-to-one correspondence between $\theta$-compatible reductions of $\mathscr{E}$ and $\nabla$-compatible reductions of $\mathscr{V}$. In this proof, we start with a $\theta$-compatible reduction $\sigma$ and construct a $\nabla$-compatible reduction $\varsigma$. The other direction can be proved in a similar way and we omit the details.
Let $(\mathscr{E},\theta)\in{\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$ be a $G$-Higgs bundle. Under the isomorphism $\mathscr{X}\cong[Y/R]$, we regard $(\mathscr{E},\theta)$ as a tuple $(E,\theta,\tau)$, where $(E,\theta)$ is a logarithmic $G$-Higgs bundle on $Y^{\prime}$ and $\tau:p_{1}^{*}(E,\theta)\cong p_{2}^{*}(E,\theta)$ is an isomorphism as in the proof of Theorem 4.2. A $\theta$-compatible reduction $\sigma:\mathscr{X}^{\prime}\rightarrow\mathscr{E}/P$ can be regarded as a $\theta$-compatible reduction $\sigma:Y^{\prime}\rightarrow E/P$ such that the following diagram commutes,
${p_{1}^{*}E_{\sigma}}$${p_{1}^{*}E}$${p_{2}^{*}E_{\sigma}}$${p_{2}^{*}E}$$\scriptstyle{p_{1}^{*}\sigma}$$\scriptstyle{\tau_{\sigma}}$$\scriptstyle{\tau}$$\scriptstyle{p_{2}^{*}\sigma}$
where $\tau_{\sigma}:p_{1}^{*}E_{\sigma}\rightarrow p_{2}^{*}E_{\sigma}$ is the isomorphism induced by $\tau$. Moreover, $p_{i}^{*}\sigma$ is compatible with the Higgs field $p_{i}^{*}\theta$. By Theorem 3.6, the triple $(E,\theta,\sigma)$ on $Y^{\prime}$ corresponds to a triple $(V,\nabla,\varsigma)$, where $(V,\nabla)$ is a nilpotent logarithmic flat $G$-bundle and $\varsigma:Y\rightarrow V/P$ is a $\nabla$-compatible reduction. Similarly, $p_{i}^{*}(E,\theta,\sigma)$ corresponds to $p_{i}^{*}(V,\nabla,\varsigma)$ on $R$. The isomorphisms
$\tau:p_{1}^{*}(E,\theta)\cong p_{2}^{*}(E,\theta),\quad\tau_{\sigma}:p_{1}^{*}(E_{\sigma},\theta_{\sigma})\cong p_{2}^{*}(E_{\sigma},\theta_{\sigma})$
@@
$F_{1}^{*}\tau:p_{1}^{*}(V,\nabla)\cong p_{2}^{*}(V,\nabla),\quad F_{1}^{*}\tau_{\sigma}:p_{1}^{*}(V_{\varsigma},\nabla_{\varsigma})\cong p_{2}^{*}(V_{\varsigma},\nabla_{\varsigma})$
by Theorem 4.2. Thus, we obtain a triple $(V,\nabla,\varsigma)$ together with an isomorphism $F_{1}^{*}\tau$ such that the diagram
[…]
@@ 5. Logahoric Higgs Torsors and Logahoric Connections
The goal of this section is to prove a logahoric version of nonabelian Hodge correspondence for $G$-bundles via exponential twisting. In §5.1, we review the construction of parahoric torsors over a smooth variety $X$. The case on curves is studied by Balaji–Seshadri [BS15] and the higher dimensional case is considered by Balaji–Pandey recently [BP22]. In §5.2, we prove a correspondence between parahoric torsors on $X$ and $G$-bundles on a root stack $\mathscr{X}$ (Proposition 5.5), and the correspondence also preserves the stability condition (Proposition 5.9). In §5.3, we give the proof of the main result (Theorem 5.17).
@@ 5.1. Parahoric Group Scheme
Recall that $G$ is a split connected reductive linear algebraic group over $k$. We fix a maximal torus $T\subseteq G$ and denote by $\mathcal{R}$ the root system. Define $X(T):={\rm Hom}(T,\mathbb{G}_{m})$ and $Y(T):={\rm Hom}(\mathbb{G}_{m},T)$. Denote by
@@
Let $X$ be a smooth variety over $k$ with a given reduced normal crossing divisor $D=\sum\limits_{i=1}^{s}D_{i}$, where $D_{i}$ are irreducible components for $1\leq i\leq s$. We also equip each $D_{i}$ with a tame weight $\alpha_{i}$, and denote by $\boldsymbol{\alpha}$ the collection of these weights. Furthermore, denote by $\boldsymbol{d}=\{d_{i},1\leq i\leq s\}$ the collection of the corresponding denominators. Recall that $\eta_{i}$ is the generic point of $D_{i}$ and $R_{i}=\mathcal{O}_{X,\eta_{i}}$ with fraction field $K_{i}$. Denote by $\widehat{R}_{i}$ and (resp. $\widehat{K}_{i}$) the completion, which is regarded as the formal neighbourhood of the generic point $\eta_{i}$. Let $G_{\alpha_{i}}(\widehat{R}_{i})$ be the parahoric group associated to a given weight $\alpha_{i}$
@@
Theorem 5.1 (Theorem 5.4 in [BP22]).
There exists a smooth affine group scheme $\mathcal{G}_{\boldsymbol{\alpha}}$ on $X$ such that
  • $\mathcal{G}_{\boldsymbol{\alpha}}|_{X\backslash D}\cong G\times X\backslash D$,
  • $\mathcal{G}_{\boldsymbol{\alpha}}|_{{\rm Spec}\,\widehat{R}_{i}}\cong\mathcal{G}_{\alpha_{i}}$.
This group scheme is unique up to isomorphism, and is called a parahoric group scheme. Moreover, there exists a vector bundle $\mathfrak{g}_{\boldsymbol{\alpha}}$ with a natural Lie algebra structure on its fiber such that
  • $\mathfrak{g}_{\boldsymbol{\alpha}}|_{X\backslash D}\cong\mathfrak{g}\times X\backslash D$,
  • $\mathfrak{g}_{\boldsymbol{\alpha}}|_{{\rm Spec}\,\widehat{R}_{i}}\cong\mathfrak{g}_{\alpha_{i}}$.
@@
There is a well-defined $\mathcal{G}_{\boldsymbol{\alpha}}$-action on $\mathfrak{g}_{\boldsymbol{\alpha}}$ induced by the adjoint action locally, and we refer the reader to [Boa11] for more details. Therefore, given a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor $\mathcal{E}$, we can define its adjoint bundle $\mathcal{E}(\mathfrak{g})$ naturally.
Example 5.3.
In this example, we give a more precise description about $\mathcal{G}_{\boldsymbol{\alpha}}$ around a formal neighborhood of a point $x\in D$. For each root $r\in\mathcal{R}$, we have a natural morphim
$u_{r}:\mathbb{G}_{a}\rightarrow G,$
and denote by $U_{r}$ the image, which is a unipotent group. Let $\widehat{\mathcal{O}}_{X,x}\cong k[[z_{1},\dots,z_{d}]]$ be a formal neighborhood of $x\in D\subseteq X$ with $\widehat{K}$ the fraction field. Suppose that $D_{i}=Z(z_{i})$ for $1\leq i\leq s$, and we assume that $s\leq d$ for convenience. Then, $\mathcal{G}_{\boldsymbol{\alpha}}|_{\widehat{\mathcal{O}}_{X,x}}$ is isomorphic to the group scheme on ${\rm Spec}\,\widehat{\mathcal{O}}_{X,x}$ determined by the following generators
$\langle T(\widehat{\mathcal{O}}),\,U_{r}((\prod_{i=1}^{s}z_{i}^{m_{r}(\alpha_{i})})\widehat{\mathcal{O}}),\,r\in\mathcal{R}\rangle.$
This group scheme $\mathcal{G}_{\boldsymbol{\alpha}}|_{\widehat{\mathcal{O}}_{X,x}}$ is called a n-parahoric group scheme [BP22, §5]. Similarly, $\mathfrak{g}_{\boldsymbol{\alpha}}|_{\widehat{\mathcal{O}}_{X,x}}$ is isomorphic to a Lie subalgebra in $\mathfrak{g}(\widehat{K})$ generated by
$\langle\mathfrak{t}(\widehat{\mathcal{O}}),\,\mathfrak{g}_{r}((\prod_{i=1}^{s}z_{i}^{m_{r}(\alpha_{i})})\widehat{\mathcal{O}}),\,r\in\mathcal{R}\rangle.$
@@
We fix the following data first:
  • $X$: smooth variety,
  • $D$: reduced normal crossing divisor with irreducible components $D_{i}$, $1\leq i\leq s$,
  • $\boldsymbol{\alpha}$: a collection of tame weights.
By Theorem 5.1, we obtain a parahoric group scheme $\mathcal{G}_{\boldsymbol{\alpha}}$ on $X$. Now we consider parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors. Given a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor, we associate it with a tuple $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\Theta_{i})$, where
@@
$\Theta_{i}:\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{i}}{\rm Spec}\,\widehat{K}_{i}\rightarrow\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,\widehat{K}_{i},$
@@
is an isomorphism for each $i$. In this case, the isomorphism $\Theta_{i}$ can be regarded as an element in $G(\widehat{K}_{i})$.
For the other direction, we start with a such a tuple $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\Theta_{i})$, where
  • $\mathcal{E}_{0}$ is a $G$-bundle on $X\backslash D$;
  • ^$\widehat{\mathcal{E}}_{i}$ is a parahoric $\mathcal{G}_{\alpha_{i}}$-torsor on ${\rm Spec}\,\widehat{R}_{i}$;
  • $\Theta_{i}:\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{i}}{\rm Spec}\,\widehat{K}_{i}\rightarrow\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,\widehat{K}_{i}$ is an isomorphism,
and we want to construct a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor on $X$. Similar to the discussion in §4.2, if the intersection $D_{i}\cap D_{j}$ is nonempty for some distinct $i,j$, such a tuple $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\Theta_{i})$ may not determine a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor. We follow the same notations as in §4.2 and introduce the compatiblility condition in a similar way. Given such a tuple $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\Theta_{i})$, if there is an isomorphism
$\Theta_{ij}:\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{ij}}{\rm Spec}\,\widehat{K}_{ij}\rightarrow\widehat{\mathcal{E}}_{j}\times_{{\rm Spec}\,\widehat{R}_{ij}}{\rm Spec}\,\widehat{K}_{ij}$
such that $(q_{0i}\circ\Theta_{i})(s_{i})=(q_{0j}\circ\Theta_{j})(s_{j})$ if and only if $(\Theta_{ij}\circ q_{ij})(s_{i})=q_{ji}(s_{j})$ for any section $s_{i}$ (resp. $s_{j}$) of ^$\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{i}}{\rm Spec}\,\widehat{K}_{i}$ (resp. ^$\widehat{\mathcal{E}}_{j}\times_{{\rm Spec}\,\widehat{R}_{j}}{\rm Spec}\,\widehat{K}_{j}$), then we say that the tuple (,^$(\mathcal{E},\widehat{\mathcal{E}}_{i},\Theta_{i})$ is compatible.
${\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,\widehat{K}_{ij}}$${\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,\widehat{K}_{i}}$${\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,\widehat{K}_{j}}$^${\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{i}}{\rm Spec}\,\widehat{K}_{i}}$^${\widehat{\mathcal{E}}_{j}\times_{{\rm Spec}\,\widehat{R}_{j}}{\rm Spec}\,\widehat{K}_{j}}$^${\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{ij}}{\rm Spec}\,\widehat{K}_{ij}}$^${\widehat{\mathcal{E}}_{j}\times_{{\rm Spec}\,\widehat{R}_{ij}}{\rm Spec}\,\widehat{K}_{ij}}$$\scriptstyle{q_{0i}}$$\scriptstyle{p_{0j}}$$\scriptstyle{q_{ij}}$$\scriptstyle{\Theta_{i}}$$\scriptstyle{q_{ji}}$$\scriptstyle{\Theta_{j}}$$\scriptstyle{\Theta_{ij}}$
As an analogue of Lemma 4.6, we have the following lemma for parahoric torsors.

### v2

@@ 4.4. Stability Condition
We remind the reader that the degree of locally free sheaves on $\mathscr{X}^{\prime}$ is defined by $H^{\prime}:=\pi^{*}_{X/k}H$, while the degree of locally free sheaves on $\mathscr{X}$ is given by $H$. It is easy to check that $\deg_{H^{\prime}}\mathscr{F}^{\prime}=p\deg_{H}\mathscr{F}$, where $\mathscr{F}$ is a locally free sheaf on $\mathscr{X}$. As discussed in Theorem 3.8, we only have to check that there is a one-to-one correspondence between $\theta$-compatible reductions of a $G$-Higgs bundle (.$(\mathscr{E}.\theta)$ on $\mathscr{X}^{\prime}$ and $\nabla$-compatible reductions of the corresponding flat $G$-bundle $(\mathscr{V},\nabla)$ on $\mathscr{X}$ given by Theorem 4.12. In this proof, we start with a $\theta$-compatible reduction $\sigma$ and construct a $\nabla$-compatible reduction $\varsigma$. The other direction can be proved in a similar way and we omit the details.
Let $(\mathscr{E},\theta)\in{\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$ be a $G$-Higgs bundle. Under the isomorphism $\mathscr{X}\cong[Y/R]$, we regard $(\mathscr{E},\theta)$ as a tuple $(E,\theta,\tau^{\prime})$, where $(E,\theta)$ is a logarithmic $G$-Higgs bundle on $Y^{\prime}$ and $\tau^{\prime}:p_{1}^{*}(E,\theta)\cong p_{2}^{*}(E,\theta)$ is an isomorphism as in the proof of Theorem 4.2. A $\theta$-compatible reduction $\sigma:\mathscr{X}^{\prime}\rightarrow\mathscr{E}/P$ can be regarded as a $\theta$-compatible reduction $\sigma:Y^{\prime}\rightarrow E/P$ such that the following diagram commutes,
${p_{1}^{*}E_{\sigma}}$${p_{1}^{*}E}$${p_{2}^{*}E_{\sigma}}$${p_{2}^{*}E}$$\scriptstyle{p_{1}^{*}\sigma}$$\scriptstyle{\tau^{\prime}_{\sigma}}$$\scriptstyle{\tau^{\prime}}$$\scriptstyle{p_{2}^{*}\sigma}$
where $\tau^{\prime}_{\sigma}:p_{1}^{*}E_{\sigma}\rightarrow p_{2}^{*}E_{\sigma}$ is the isomorphism induced by $\tau^{\prime}$. Moreover, $p_{i}^{*}\sigma$ is compatible with the Higgs field $p_{i}^{*}\theta$. By Theorem 3.8, the triple $(E,\theta,\sigma)$ on $Y^{\prime}$ corresponds to a triple $(V,\nabla,\varsigma)$, where $(V,\nabla)$ is a nilpotent logarithmic flat $G$-bundle and $\varsigma:Y\rightarrow V/P$ is a $\nabla$-compatible reduction. Similarly, $p_{i}^{*}(E,\theta,\sigma)$ corresponds to $p_{i}^{*}(V,\nabla,\varsigma)$ on $R$. The isomorphisms
$\tau^{\prime}:p_{1}^{*}(E,\theta)\cong p_{2}^{*}(E,\theta),\quad\tau^{\prime}_{\sigma}:p_{1}^{*}(E_{\sigma},\theta_{\sigma})\cong p_{2}^{*}(E_{\sigma},\theta_{\sigma})$
@@
$\tau:p_{1}^{*}(V,\nabla)\cong p_{2}^{*}(V,\nabla),\quad\tau_{\sigma}:p_{1}^{*}(V_{\varsigma},\nabla_{\varsigma})\cong p_{2}^{*}(V_{\varsigma},\nabla_{\varsigma})$
by Theorem 4.2. Thus, we obtain a triple $(V,\nabla,\varsigma)$ together with an isomorphism $\tau$ such that the diagram
${p_{1}^{*}V_{\varsigma}}$${p_{1}^{*}V}$${p_{2}^{*}V_{\varsigma}}$${p_{2}^{*}V}$$\scriptstyle{p_{1}^{*}\varsigma}$$\scriptstyle{\tau_{\sigma}}$$\scriptstyle{\tau}$$\scriptstyle{p_{2}^{*}\varsigma}$
@@ 5. Logahoric Higgs Torsors and Logahoric Connections
The goal of this section is to prove a logahoric version of the nonabelian Hodge correspondence for $G$-bundles via exponential twisting. We first studied parahoric group schemes over a smooth variety $X$ (with a given reduced normal crossing divisor $D$), which is inspired by Balaji–Pandey’s work [BP22, Theorem 5.4]. Then we prove a correspondence between parahoric torsors on $X$ and $G$-bundles on a root stack $\mathscr{X}$ (Proposition 5.5), which also preserves the stability condition (Proposition 5.10). Based on the above correspondence and results given on root stacks (Theorem 4.12 and 4.15), we give the proof of the main result (Theorem 5.18).
@@ 5.1. Parahoric Group Scheme
We fix a maximal torus $T\subseteq G$ and denote by $\mathcal{R}$ the root system. Define $X(T):={\rm Hom}(T,\mathbb{G}_{m})$ and $Y(T):={\rm Hom}(\mathbb{G}_{m},T)$. Denote by
@@
Let $X$ be a smooth variety over $k$ with a given reduced normal crossing divisor $D=\sum\limits_{i=1}^{s}D_{i}$, where $D_{i}$ are irreducible components for $1\leq i\leq s$. We equip each $D_{i}$ with a tame weight $\alpha_{i}$, and denote by $\boldsymbol{\alpha}$ the collection of these weights. Furthermore, denote by $\boldsymbol{d}=\{d_{i},1\leq i\leq s\}$ the collection of the corresponding denominators. Recall that $\eta_{i}$ is the generic point of $D_{i}$ and $R_{i}=\mathcal{O}_{X,\eta_{i}}$ with fraction field $K_{i}$. Denote by $\widehat{R}_{i}$ and (resp. $\widehat{K}_{i}$) the completion, which is regarded as the formal neighbourhood of the generic point $\eta_{i}$. Let $G_{\alpha_{i}}(\widehat{R}_{i})$ be the parahoric group associated to a given weight $\alpha_{i}$
@@
Definition 5.1.
A smooth group scheme $\mathcal{G}$ on $X$ is parahoric if
$\mathcal{G}\times_{X}X\backslash D\cong G\times X\backslash D$
and
$\mathcal{G}\times_{X}{\rm Spec}\,\widehat{R}_{i}\cong\mathcal{G}_{\alpha_{i}}$
for some tame weights $\alpha_{i}$ for $1\leq i\leq s$.
The idea of the above definition comes from a recent work by Balaji and Pandey [BP22], especially [BP22, Theorem 5.4], where the authors constructed parahoric group schemes on higher dimensional varieties for simply connected and almost simple algebraic group $G$. In the following, we will show that parahoric group schemes in the above sense exist.
Let $\mathscr{X}:=\mathscr{X}_{(X,D,\boldsymbol{d})}$ be a root stack with coarse moduli space $X$ studied in §4.2. Let $\mathscr{F}$ be a $G$-bundle on $\mathscr{X}$. Consider the stack of automorphisms ${\rm Aut}_{\mathscr{X}}(\mathscr{F},\mathscr{F})$. We define a functor
$(\pi_{*}{\rm Aut}_{\mathscr{X}}(\mathscr{F},\mathscr{F}))(T):={\rm Hom}_{\mathscr{X}}(T\times_{X}\mathscr{X},{\rm Aut}_{\mathscr{X}}(\mathscr{F},\mathscr{F}))$
for every $X$-scheme $T$.
Lemma 5.2.
Let $\mathscr{F}$ be a $G$-bundle of type $\boldsymbol{\rho}$ on $\mathscr{X}$ such that $\mathscr{F}|_{X\backslash D}$ is a trivial $G$-bundle. The functor $(\pi_{*}{\rm Aut}_{\mathscr{X}}(\mathscr{F},\mathscr{F}))$ is representable by a parahoric group scheme $\mathcal{G}$ on $X$ such that $\mathcal{G}\times_{X}{\rm Spec}\,\widehat{R}_{i}\cong\mathcal{G}_{\alpha_{i}}$ for $1\leq i\leq s$, where the tame weight $\alpha_{i}$ is determined by $\rho_{i}$.
Proof.
We prove this lemma under the isomorphism $\mathscr{X}\cong[Y/\Gamma]$ and regard $\mathscr{F}$ as a $(\Gamma,G)$-bundle on $Y$ (Remark 4.4). Thus the functor $\pi_{*}{\rm Aut}_{\mathscr{X}}(\mathscr{F},\mathscr{F})$ is equivalent to $(\pi_{*}{\rm Aut}_{Y}(\mathscr{F},\mathscr{F}))^{\Gamma}$. Following [BLR90, §7.6] (see also [BS15, §4] and [Dam24]), the functor $(\pi_{*}{\rm Aut}_{Y}(\mathscr{F},\mathscr{F}))^{\Gamma}$ is representable by a smooth group scheme $\mathcal{G}$ on $X$.
Now we will show that the group scheme $\mathcal{G}$ is parahoric. Note that “parahoric” is a local condition, and thus it is enough to work on $[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ for $1\leq i\leq s$ (see notations in §4.2). Given a representation $\rho_{i}:\mu_{d_{i}}\rightarrow T$, we can associate it with a tame weight $\alpha_{i}$ [BS15, Lemma 2.2.8]. Then, by [BS15, Theorem 2.3.1], we have
$\mathcal{G}\times_{X}{\rm Spec}\,\widehat{R}_{i}=\pi_{*}{\rm Aut}_{{\rm Spec}\,\widehat{\widetilde{R}}_{i}}(\mathscr{F}|_{{\rm Spec}\,\widehat{\widetilde{R}}_{i}},\mathscr{F}|_{{\rm Spec}\,\widehat{\widetilde{R}}_{i}})^{\Gamma}\cong\mathcal{G}_{\alpha_{i}}.$
Therefore, the group scheme $\mathcal{G}$ on $X$ is parahoric. ∎
Notation 5.3.
Lemma 5.2 gives a way to construct parahoric group schemes over higher dimensional varieties, which depends on a $G$-bundle on the root stack. Thus, it is meaningful to use the notation $\mathcal{G}_{\mathscr{F},\boldsymbol{\alpha}}$ for the parahoric group scheme constructed in the lemma, where $\mathscr{F}$ is the given $G$-bundle on $\mathscr{X}$ and $\boldsymbol{\alpha}$ is the collection of tame weights determined by the type of $\mathscr{F}$. However, if there is no ambiguity, we still use the notation $\mathcal{G}_{\boldsymbol{\alpha}}:=\mathcal{G}_{\mathscr{F},\boldsymbol{\alpha}}$ for simplicity.
@@
Balaji and Pandey worked on simply connected and almost simple algebraic group and constructed parahoric group schemes on smooth projective varieties (with reduced normal crossing divisors) based on the existence of a Lie algebra bundle [BP22, Theorem 5.4]. In our setup, Lemma 5.2 provides another approach to construct parahoric group schemes when $G$ is reductive. However, we do not know whether all such parahoric group schemes (Definition 5.1) can be constructed in this way. Nevertheless, if the reduced normal crossing divisor is smooth, all parahoric group schemes (Definition 5.1) can be constructed in this way because this case is exactly the same as that on curves and we can use transition functions to give the construction (see [BS15, §3] or Remark 4.7).
@@
We follow the same setup as in §5.1. Let $\mathscr{F}$ be a $G$-bundle of type $\boldsymbol{\rho}$ on $\mathscr{X}$ such that $\mathscr{F}|_{X\backslash D}$ is trivial, and denote by $\mathcal{G}_{\boldsymbol{\alpha}}$ the corresponding parahoric group scheme given by Lemma 5.2.
Proposition 5.5.
The category ${\rm Bun}(X,\mathcal{G}_{\boldsymbol{\alpha}})$ of parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors on $X$ is equivalent to the category ${\rm Bun}(\mathscr{X},G,\boldsymbol{\rho})$ of $G$-bundles of type $\boldsymbol{\rho}$ on $\mathscr{X}$.
Proof.
For convenience, we regard $\mathscr{X}$ as a global quotient $[Y/\Gamma]$, and then the proposition follows directly from [Dam24, Theorem 11]. The equivalence is given by
$\begin{split}\beta_{\mathscr{F}}:&{\rm Bun}(\mathscr{X},G,\boldsymbol{\rho})\rightarrow{\rm Bun}(X,\mathcal{G}_{\boldsymbol{\alpha}})\\ &\mathscr{F}^{\prime}\mapsto\pi_{*}{\rm Aut}_{\mathscr{X}}(\mathscr{F},\mathscr{F}^{\prime})\end{split}$
and
$\begin{split}\eta_{\mathscr{F}}:&{\rm Bun}(X,\mathcal{G}_{\boldsymbol{\alpha}})\rightarrow{\rm Bun}(\mathscr{X},G,\boldsymbol{\rho})\\ &\mathscr{E}\mapsto\pi^{*}\mathscr{E}\times_{\pi^{*}\mathcal{G}_{\boldsymbol{\alpha}}}\mathscr{F}.\end{split}$
Similar to the local discussion for $G$-bundles on Deligne–Mumford stacks in Remark 4.7, we consider the local picture of parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors. Given a parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor $\mathcal{E}$, we associate it with a tuple $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\widehat{\Theta}^{\prime}_{i})$, where
@@
$\widehat{\Theta}^{\prime}_{i}:\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,\widehat{K}_{i}\rightarrow\widehat{\mathcal{E}}_{i}\times_{{\rm Spec}\,\widehat{R}_{i}}{\rm Spec}\,\widehat{K}_{i}$
@@
is an isomorphism for each $i$. In this case, the isomorphism $\widehat{\Theta}^{\prime}_{i}$ can be regarded as an element in $G(\widehat{K}_{i})$.
Remark 5.6.
In this remark, we briefly explain Proposition 5.5 in the viewpoint of changing transition functions. We first recall some notations. Let $R_{i}$ be the discrete valuation rings given by the generic point $\eta_{i}$ of $D_{i}$ with fraction field $K_{i}$. We have
Let $\mathscr{E}$ be a $G$-bundle of type $\boldsymbol{\rho}$ on $\mathscr{X}$ and denote by $\mathcal{E}$ the corresponding parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor on $X$. Denote by $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\widehat{\Theta}_{i})$ (resp. $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\widehat{\Theta}^{\prime}_{i})$) the tuple associated to $\mathscr{E}$ (resp. $\mathcal{E}$). We will explain the relation between $\widehat{\Theta}_{i}$ and $\widehat{\Theta}^{\prime}_{i}$ as follows. Recall that
$\widehat{\Theta}_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\widehat{\Theta}_{i}(w_{i}).$
By [BS15, (2.2.9.4)], we can choose an element $\Delta_{i}\in T(\widehat{\widetilde{R}}_{i})$ such that
$\Delta_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\Delta_{i}(w_{i}),$
and define new isomorphisms
[…]
It is easy to check that $\Delta^{-1}_{i}\widehat{\Theta}_{i}$ is $\mu_{d_{i}}$-invariant, i.e.
$(\Delta^{-1}_{i}\widehat{\Theta}_{i})(\gamma_{i}\cdot w_{i})=(\Delta^{-1}_{i}\widehat{\Theta}_{i})(w_{i}).$
Thus, $\Delta^{-1}_{i}\widehat{\Theta}_{i}$ descents to a well-defined element in $G(\widehat{K}_{i})$ under the substitution $z_{i}=w_{i}^{d_{i}}$, which is exactly $\widehat{\Theta}^{\prime}_{i}$ (up to conjugation). Moreover, there is a one-to-one correspondence between elements in $G(\widehat{\widetilde{K}}_{i})^{\rho_{i}}$ and elements in $G(\widehat{K}_{i})$ in the above sense.
At the end of this remark, we briefly explain the relation between tuples $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\widehat{\Theta}^{\prime}_{i})$ (resp. $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\widehat{\Theta}_{i})$) and parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsors $\mathcal{E}$ on $X$ (resp. $G$-bundles on $\mathscr{X}$). We take parahoric torsors as an example and the approach given here is also included in [BP22, first and second paragraph of the proof of Theorem 5.4]. We start with a tuple $(\mathcal{E}_{0},\widehat{\mathcal{E}}_{i},\widehat{\Theta}^{\prime}_{i})$. By [BLR90, §6], the tuple induces one $(\mathcal{E}_{0},\mathcal{E}_{i},\Theta^{\prime}_{i})$, where $\mathcal{E}_{i}$ is a scheme of finite type on ${\rm Spec}\,R_{i}$ and
$\Theta^{\prime}_{i}:\mathcal{E}_{0}\times_{X\backslash D}{\rm Spec}\,K_{i}\rightarrow\mathcal{E}_{i}\times_{{\rm Spec}\,R_{i}}{\rm Spec}\,K_{i}$
is an isomorphism. Therefore, the tuple $(\mathcal{E}_{0},\mathcal{E}_{i},\Theta^{\prime}_{i})$ gives a scheme $\mathcal{E}^{\rm big}$ over an open subset $X^{\rm big}$ of $X$ such that ×$\mathcal{E}\times_{X}X^{\rm big}\cong\mathcal{E}^{\rm big}$ and ${\rm codim}(X\backslash X^{\rm big})\geq 2$. Moreover, $\mathcal{E}^{\rm big}$ is a parahoric $\mathcal{G}^{\rm big}_{\boldsymbol{\alpha}}$-torsor, where $\mathcal{G}^{\rm big}_{\boldsymbol{\alpha}}:=\mathcal{G}_{\boldsymbol{\alpha}}\times_{X}X^{\rm big}$.
Remark 5.7.
In this remark, we mention a property, which will be used to study how Higgs fields and connections change later. Recall that we define
$G(\widehat{\widetilde{R}}_{i})^{\rho_{i}}:=\{\widehat{\Psi}_{i}\in G(\widehat{\widetilde{R}}_{i})\,|\,\rho_{i}(\gamma_{i})\widehat{\Psi}_{i}(w_{i})\rho_{i}(\gamma_{i})^{-1}=\Psi_{i}(\gamma_{i}\cdot w_{i})\}.$
in Remark 4.7. With a similar approach as [BS15, Theorem 2.3.1], we have
More precisely, given an isomorphism $\widehat{\Psi}_{i}:\mathscr{E}_{i}\rightarrow\mathscr{E}_{i}$, we regard it as an element in $G(\widehat{\widetilde{R}}_{i})^{\rho_{i}}$, i.e.
$\rho(\gamma_{i})\widehat{\Psi}_{i}(w_{i})\rho(\gamma_{i})^{-1}=\widehat{\Psi}_{i}(\gamma_{i}\cdot w_{i}).$
Define $\widehat{\Psi}^{\prime}_{i}=\Delta_{i}^{-1}\widehat{\Psi}_{i}\Delta_{i}$, which is $\mu_{d_{i}}$-invariant, i.e.
$\widehat{\Psi}^{\prime}_{i}(\gamma_{i}\cdot w_{i})=\widehat{\Psi}^{\prime}_{i}(w_{i}).$
Thus, $\widehat{\Psi}^{\prime}_{i}(w_{i})$ descents to a well-defined element in $G(\widehat{K}_{i})$ and it is easy to check that it is in $G_{\alpha_{i}}(\widehat{R}_{i})$. Indeed, conjugating by $\Delta_{i}$ and substituting $z_{i}=w_{i}^{d_{i}}$ induces the desired isomorphism.
Now we suppose that $X$ is projective, and then $Y$ given in Remark 4.4 can be required to be projective by Kawamata covering lemma. Let $\eta$ (resp. $\widetilde{\eta}$) be the generic point of $X$ (resp. $Y$). Let $\mathcal{E}$ be parahoric $\mathcal{G}_{\boldsymbol{\alpha}}$-torsor on $X$. Let $\mathcal{P}$ be a maximal parabolic subgroup of the generic fiber $(\mathcal{G}_{\boldsymbol{\alpha}})_{\eta}$ of $\mathcal{G}_{\boldsymbol{\alpha}}$. Given a generic reduction $\varsigma_{\eta}:X_{\eta}\rightarrow\mathcal{E}_{\eta}/\mathcal{P}$ and a character $\kappa:\mathcal{P}\rightarrow\mathbb{G}_{m,\eta}$, we obtain a line bundle $L(\varsigma_{\eta},\kappa)$ on $X\backslash D$.